# Structured Reinforcement Learning for Media Streaming at the Wireless Edge

### v1

@@ 4.1. Convergence of NPG for Threshold Parameterization
(11) $\lambda_{t+1}$ $=\mathcal{P}_{\Lambda}(\lambda_{t}+\eta_{2}\sum_{n=1}^{N}\nabla_{\lambda}J^{\pi_{\theta_{n,t}}}(\rho_{n};\lambda_{t}))$
@@
Note that Algorithm (4.1) is a straightforward extension of the standard primal dual approach to solve the saddle point problem (4). The primal variables here are the threshold parameters of each client $n\in[N]$. At each iteration $t+1$, the threshold parameters of client $n$ follow an NPG update independent of the others. Intuitively, the update makes sense as it is known from Theorem (1), it is sufficient to look at client policies which are independent of other but the respective value functions are coupled only through the dual variable $\lambda_{t}$ from the previous iteration $t$. The dual variable update at each iteration, however, uses the sum of gradients from each client $n$. It is easy to see that the our algorithm reduces to the standard primal dual method when $N=1$.
The following lemma shows that the soft threshold parameterization has a rather simple update rule under the natural policy gradient method. Denoting actions $1$,$2$ by $a_{1}$ and $a_{2}$, respectively.
@@
For the soft threshold policy parameterization, the following are equivalent:
@@
It is easy to show that for the soft threshold policy class, we have
(12) $\frac{\partial{\log\pi_{\theta}(s,a)}}{\partial{\theta_{s^{\prime}}}}$
@@
and therefore,
(14) $\pi_{\theta}(s,a_{1})\nabla_{\theta}\log\pi_{\theta}(s,a_{1})+\pi_{\theta}(s,a_{2})\nabla_{\theta}\log\pi_{\theta}(s,a_{2})=0.$
@@
(15) $\nabla_{\theta}\log\pi_{\theta}(s,a_{1})-\nabla_{\theta}\log\pi_{\theta}(s,a_{2})=e_{s}.$
where $e_{s}\in\mathbb{R}^{\absolutevalue{\mathcal{S}}}$ such that the $i$-th entry is $1$ if $i=s$ and 00 otherwise. We now use the fact that the direction of primal update in  (4.1), parallels the solution of the following minimization problem.
@@
That is, the optimal solution to the above minimization problem $w^{*}=(1-\gamma)F(\theta_{n,t})^{\dagger}\nabla_{\theta}J^{\pi_{\theta_{n,t}}}(\rho_{n};\lambda_{t})$. On the other hand,
(17) $\mathbb{E}_{s\sim d_{\rho_{n}}^{\pi_{\theta_{n,t}}},a\sim\pi_{\theta_{n,t}}(s,\cdot)}[(A^{\pi_{\theta_{n,t}}}_{\lambda_{t}}(s,a)-w\cdot\nabla_{\theta}\log\pi_{\theta_{n,t}}(s,a))^{2}]$
@@
*$*\pi_{\theta_{n,t}}(s,a_{1})w\cdot\nabla_{\theta}\log\pi_{\theta_{n,t}}(s,a_{1}))$
(18) $+\pi_{\theta_{n,t}}(s,a_{1})w\cdot\nabla_{\theta}\log\pi_{\theta_{n,t}}(s,a_{1})(w\cdot e_{s})]$
@@
(19) $(w\cdot e_{s}-2(A^{\pi_{\theta_{n,t}}}_{\lambda_{t}}(s,a_{1})-A^{\pi_{\theta_{n,t}}}_{\lambda_{t}}(s,a_{2})))]$
The first equality follows from expanding the quadratic expansion and (14), and the last equality from (15). It is easy to see that the optimal solution $w^{*}$ is also the optimal minimizer for the second term in the last equality. Using the fact that the stochastic policy $\pi_{\theta_{n,t}}(s,a)>0$ for all state-action pairs, we have $w^{*}\cdot e_{s}=(A^{\pi_{\theta_{n,t}}}_{\lambda_{t}}(s,a_{1})-A^{\pi_{\theta_{n,t}}}_{\lambda_{t}}(s,a_{2}))$
An immediate consequence of lemma (4) is the following form for the policy updates.
@@
(21) $\pi_{\theta_{n,t+1}}(s,a)$ $=\pi_{\theta_{n,t}}(s,a)\frac{\exp(\frac{-\eta_{1}}{1-\gamma}A^{\pi_{\theta_{n,t}}}_{\lambda_{t}}(s,a))}{Z_{t}(s)}.$
[…]
The proofs of Corollary 1 is given in Appendix .3
Evaluating the gradient with respect to the dual variable in the dual update, the equivalent primal dual method to algorithm (4.1) is given by,
@@
(23) $\lambda_{t+1}=\mathcal{P}_{\Lambda}(\lambda_{t}+\eta_{2}(\sum_{n=1}^{N}J_{g}^{\pi_{\theta_{n,t}}}(\rho_{n})-\bar{K}))$
We have the following convergence guarantee for algorithm. (4.1).
@@
(24) $\frac{1}{T}\sum_{t=0}^{T-1}\sum_{n=1}^{N}J_{c}^{\pi_{\theta_{n,t}}}(\rho_{n})-J_{c}^{\pi_{n}^{*}}(\rho_{n})$ $\leq\frac{4N}{(1-\gamma)^{2}\sqrt{T}}.$
(25) $(\frac{1}{T}\sum_{t=0}^{T-1}\sum_{n=1}^{N}J_{g}^{\pi_{\theta_{n,t}}}(\rho)-\bar{K})^{+}$ $\leq\frac{(2/\xi+4\xi)N^{2}}{(1-\gamma)^{2}\sqrt{T}}.$
The proof technique we follow is similar to that of  (Ding et al., 2020a). We first show that the NPG method has a simple parameter update for the threshold policy class. We then prove that the average cost function generated by the iterates of the algorithm converges to the global optimal value, and bound the constraint violation. The details are provided in Appendix .4.
4.1.1. Algorithm Design
Algorithm 1 summarizes the primal dual method 4.1 consistent with the standard actor-critic implementations of reinforcement learning algorithms. We iteratively update the value function (and hence the advantage function), the policy parameter, and the lagrange multiplier. At each time step, we run one actor-critic update for each client. In particular, for each client $n\in N$, we sample an action according to the current policy of the client to update the value function of the current state $J_{n,t+1}(s_{n}(t))$ from which the advantage function for each action $a\in\mathcal{A}$, $A_{n,t}(s_{n}(t),a)$ is computed. The threshold parameter for the current state is updated according to  (4.1). The value functions and parameters of all but the current state remain unchanged. After each client updates its value function and its policy, the experienced actions are retrieved, and the lagrange multiplier is updated using the constraint formulation.
4.1.2. Note
@@
Figure 2. Training in simulation
Figure 3. QoE in simulation
Figure 4. Clients assigned to high priority service
@@ 5.2. Training and Evaluation
Figure 4 shows the evaluation of the algorithms on a joint system during training (averaged over $5$ random seeds). We observe that decentralized algorithms converge over four times faster than centralized algorithms as they exploit the structure of the environment. The performance difference between the trained centralized and decentralized algorithms is negligible as seen in Figure 4 (averaged over $100$ runs). The performance of DC-T is on par with the best performing algorithms, which confirms our hypothesis that the optimal policy is indeed a threshold policy. The index policy, I that is hard-constrained version of DC-T shows similarly high performance. Figure 4 (averaged over $100$ runs) shows the number of clients in the high priority queue during evaluation of the soft constrained algorithms is near 2, and so confirms our choice of $\lambda$. Finally, Figure 5 shows the efficacy of the DC-T algorithm in deployment in terms of inference times. We can see that the PPO algorithms take roughly $70\mu s$, while the DC-T algorithm takes less than $20\mu s$ for inferring the decisions for the $6$ client system.
@@ 6. Real-World Evaluation
Figure 8. Comparison of average buffer state
Figure 9. Comparison of average QoE with varying number clients
Figure 10. Comparison of QoE CDF with varying number of clients
Figure 12. Comparison of average QoE in a poor channel
Figure 13. Comparison of QoE CDF in a poor channel
Figure 6. Comparison of average QoE
Figure 7. Comparison of QoE CDF
Figure 8. Comparison of average buffer state
@@ C. Does the index approach generalize to variable channels?
Wireless channels are subject to variability due to mobility and obstruction, and the question arises as to whether the RL approach needs to be trained separately over many channel realizations? Since indexing simply orders clients based on their states, we wish to determine if it can be used unchanged even when the channel quality is low. Hence, we introduce packet losses and delays using a network emulator to create channels with disturbances. We first group clients with similar channels into clusters and divide the total available resources in a proportionally fair manner across clusters. We then apply our service differentiation policies across members of each cluster. We show results for a cluster seeing 10% packet loss and 20 ms delay, and, as expected, QoE drops as seen in Figure 14. However, the overall order still holds as Vanilla and Greedy Buffer have QoEs over 20% worse than Index. Further, the client QoE CDF and average buffer state indicate same trend as seen in 14 and 14, with Indexing still performing significantly better than the base approaches.

### v2

@@ 4.1. Convergence of NPG for Threshold Parametrization
(9) $\lambda_{t+1}$ $=\mathcal{P}_{\Lambda}\left(\lambda_{t}+\eta_{2}\sum_{n=1}^{N}\nabla_{\lambda}J^{\pi_{\theta_{n,t}}}(\rho_{n};\lambda_{t})\right),$
@@
Note that Algorithm (4.1) is a straightforward extension of the standard primal-dual approach to solve the saddle point problem (4). The primal variables here are the threshold parameters of each client $n\in[N]$. At each iteration $t+1$, the threshold parameters of client $n$ follow an NPG update independent of the others. Intuitively, the update makes sense as it is known from Theorem 1 that it is sufficient to consider client policies that are independent of each other, but the respective value functions are coupled only through the dual variable $\lambda_{t}$ from the previous iteration $t$. The dual variable update at each iteration, however, uses the sum of gradients from each client $n$. It is easy to see that the our approach reduces to the standard primal-dual method when $N=1$.
The following lemma shows that the soft-threshold parametrization has a rather simple update rule under the natural policy gradient method. Denoting actions $1$, $2$ by $a_{1}$ and $a_{2}$, respectively.
@@
For the soft-threshold policy parametrization, the following are equivalent:
@@
    $\theta_{n,{t+1}}(s)=\theta_{n,t}(s)-\frac{\eta_{1}}{(1-\gamma)}[A^{\pi_{\theta_{n,t}}}_{\lambda_{t}}(s,a_{1})-A^{\pi_{\theta_{n,t}}}_{\lambda_{t}}(s,a_{2})].$
@@
It is easy to show that for the soft-threshold policy class,
$\frac{\partial{\log\pi_{\theta}(s,a)}}{\partial{\theta_{s^{\prime}}}}$
@@
Therefore,
(11) $\pi_{\theta}(s,a_{1})\nabla_{\theta}\log\pi_{\theta}(s,a_{1})+\pi_{\theta}(s,a_{2})\nabla_{\theta}\log\pi_{\theta}(s,a_{2})=0,$
@@
(12) $\nabla_{\theta}\log\pi_{\theta}(s,a_{1})-\nabla_{\theta}\log\pi_{\theta}(s,a_{2})=e_{s},$
where $e_{s}\in\mathbb{R}^{\absolutevalue{\mathcal{S}}}$ such that the $i$-th entry is $1$ if $i=s$ and 00 otherwise.
We now consider the following quadratic minimization problem:
@@
The optimal solution to the above minimization problem is given by $w^{*}=(1-\gamma)F(\theta_{n,t})^{\dagger}\nabla_{\theta}J^{\pi_{\theta_{n,t}}}(\rho_{n};\lambda_{t})$.
Notice that the direction of primal update in equations (4.1) parallels this solution. Using this fact, we have,
$\mathbb{E}_{s\sim d_{\rho_{n}}^{\pi_{\theta_{n,t}}},a\sim\pi_{\theta_{n,t}}(s,\cdot)}[(A^{\pi_{\theta_{n,t}}}_{\lambda_{t}}(s,a)-w\cdot\nabla_{\theta}\log\pi_{\theta_{n,t}}(s,a))^{2}]$
@@
$*\pi_{\theta_{n,t}}(s,a_{1})w\cdot\nabla_{\theta}\log\pi_{\theta_{n,t}}(s,a_{1}))$
$+\pi_{\theta_{n,t}}(s,a_{1})w\cdot\nabla_{\theta}\log\pi_{\theta_{n,t}}(s,a_{1})(w\cdot e_{s})]$
@@
$(w\cdot e_{s}-2(A^{\pi_{\theta_{n,t}}}_{\lambda_{t}}(s,a_{1})-A^{\pi_{\theta_{n,t}}}_{\lambda_{t}}(s,a_{2})))]$
The first equality follows from expanding the quadratic expansion and (11), and the last equality from (12). It is easy to see that the optimal solution $w^{*}$ is also the optimal minimizer for the second term in the last equality. Using the fact that the stochastic policy $\pi_{\theta_{n,t}}(s,a)>0$ for all state-action pairs, we have $w^{*}\cdot e_{s}=(A^{\pi_{\theta_{n,t}}}_{\lambda_{t}}(s,a_{1})-A^{\pi_{\theta_{n,t}}}_{\lambda_{t}}(s,a_{2})).$
An immediate consequence of Lemma 4 is the following form for the policy updates.
@@
(14) $\pi_{\theta_{n,t+1}}(s,a)$ $=\pi_{\theta_{n,t}}(s,a)\frac{\exp(\frac{-\eta_{1}}{1-\gamma}A^{\pi_{\theta_{n,t}}}_{\lambda_{t}}(s,a))}{Z_{t}(s)},$
(15) $\text{where}\leavevmode\nobreak\ \leavevmode\nobreak\ Z_{t}(s)$ $\triangleq\sum_{a\in\mathcal{A}}\pi_{\theta_{n,t}}(s,a)\exp(\frac{-\eta_{1}}{1-\gamma}A^{\pi_{\theta_{n,t}}}_{\lambda_{t}}(s,a)).$
The proof of Corollary 1 is given in Appendix  A.3.
Evaluating the gradient with respect to the dual variable in the dual update, the equivalent primal-dual method to algorithm (4.1) is given by
@@
(16) $\lambda_{t+1}=\mathcal{P}_{\Lambda}\left(\lambda_{t}+\eta_{2}\left(\sum_{n=1}^{N}J_{g}^{\pi_{\theta_{n,t}}}(\rho_{n})-\bar{K}\right)\right).$
We have the following convergence guarantee for algorithm (4.1).
@@
(17) $\frac{1}{T}\sum_{t=0}^{T-1}\sum_{n=1}^{N}(J_{c}^{\pi_{\theta_{n,t}}}(\rho_{n})-J_{c}^{\pi_{n}^{*}}(\rho_{n}))$ $\leq\frac{4N}{(1-\gamma)^{2}\sqrt{T}}.$
(18) $\left(\frac{1}{T}\sum_{t=0}^{T-1}\sum_{n=1}^{N}J_{g}^{\pi_{\theta_{n,t}}}(\rho)-\bar{K}\right)^{+}$ $\leq\frac{(2/\xi+4\xi)N^{2}}{(1-\gamma)^{2}\sqrt{T}}.$
The proof sketch is as follows. We first show that the NPG method has a simple parameter update for the threshold policy class. We then prove that the average cost function generated by the iterates of the algorithm converges to the global optimal value, and bound the constraint violation. The details are provided in appendix.
Note that the quadratic scaling of constraint error with the number of users $N$ arises due the resources being fixed. It is easy to show that if the ability to support high-quality service, $\bar{K}$ scales linearly with $N$, the constraint error will also only be linear in $N.$
Remark: The notion of soft-threshold parametrization coupled with natural policy gradient is motivated by constrained RL using soft-max parametrization with natural policy gradient (Ding et al., 2020). However, the convergence analysis is not directly transferable from one to the other, since the structure of Fischer-information matrix induced by the soft-max policy is different from that of the soft-threshold policy, which translates to different primal dual updates in equation (4.1). Nevertheless, the overall flow of the proof is adaptable to the soft-threshold case, and we obtain a similar rate of convergence.
Algorithm Design
Algorithm 1 summarizes our approach (4.1) consistent with the standard actor-critic implementations of RL algorithms. We iteratively update the value function (and hence the advantage function), the policy parameter, and the Lagrange multiplier. At each time step, we run one actor-critic update for each client. In particular, for each client $n\in N$, we sample an action according to the current policy of the client to update the value function of the current state $J_{n,t+1}(s_{n}(t))$ from which the advantage function for each action $a\in\mathcal{A}$, $A_{n,t}(s_{n}(t),a)$ is computed. The threshold parameter for the current state is updated according to (4.1). The value functions and parameters of all but the current state remain unchanged. After each client updates its value function and its policy, the experienced actions are retrieved, and the Lagrange multiplier is updated using the constraint formulation.
Note
@@
Figure 4. Clients in high priority service
@@ 5.2. Training and Evaluation
Figure 4 shows the evaluation of the algorithms on a joint system during training (averaged over $5$ random seeds). We observe that decentralized algorithms converge over four times faster than centralized algorithms as they exploit the structure of the environment. The performance difference between the trained centralized and decentralized algorithms is negligible as seen in Figure 4 (averaged over $100$ runs). The performance of DC-T is on par with the best performing algorithms, which confirms our hypothesis that the optimal policy is indeed a threshold policy. The index policy, I that is hard-constrained version of DC-T shows similarly high performance. Figure 4 (averaged over $100$ runs) shows the number of clients in the high priority queue during evaluation of the soft constrained algorithms is near 2, and so confirms our choice of $\lambda$. Finally, Figure 5 shows the efficacy of the DC-T algorithm in deployment in terms of inference times. We see that the PPO algorithms take roughly 50 to 60 $\mu$s, while the DC-T algorithm takes only about 10 to 15 $\mu$s for inferring the decisions for the $6$ client system.
@@ 6. Real-World Evaluation
Figure 8. Comparison of average buffer
@@ C. Does the index approach generalize to variable channels?
Wireless channels are subject to variability due to mobility and obstruction, and the question arises as to whether the RL approach needs to be trained separately over many channel realizations? Since indexing simply orders clients based on their states, we wish to determine if it can be used unchanged even when the channel quality is low. Hence, we introduce packet losses and delays using a network emulator to create channels with disturbances. We first group clients with similar channels into clusters and divide the total available resources in a proportionally fair manner across clusters. We then apply our service differentiation policies across members of each cluster. We show results for a cluster seeing 10% packet loss and 20 ms delay, and, as expected, QoE drops as seen in Figure 14. However, the overall order still holds as Vanilla and Greedy Buffer have QoEs over 20% worse than Index. Further, the client QoE CDF and average buffer state indicate same trend as seen in 14 and 14, with Indexing still performing significantly better than the others.